# Post-hoc Reward Calibration: A Case Study on Length Bias

### v1

@@ 6 Conclusion
This paper introduces a post-hoc reward calibration method to mitigate biases in RMs. Our approach effectively reduces the impact of prevalent biases, such as length bias, and has been validated across multiple experimental settings. The results consistently show substantial performance gains. Other empirical findings confirm that our method satisfies several critical desiderata for optimal calibration: (1) Our approach consistently mitigates both length and markdown biases in RMs, and can successfully calibrate LLM-as-Judge rewards, underscoring its generality. (2) The method requires no additional data annotation or RM retraining and is computationally efficient. (3) The calibrated rewards exhibit weak correlations with the characteristics of interest, thereby preventing reward hacking and improving the overall quality of the reward signals. (4) For non-biased RMs, our method minimally alters the reward behaviour, whereas it substantially improves models with stronger biases, making the calibration impact proportional to the bias level. (5) The method is robust to various hyperparameter choices and can be controlled smoothly by introducing a calibration constant before LWR prediction, enhancing its practical usability. (6) The approach is not constrained by dataset size, as even a few hundred data points are sufficient to achieve effective calibration. In conclusion, our approach offers a scalable, efficient, and generalisable solution for addressing biases in RMs, ensuring better alignment between LLMs and human preferences.
Acknowledgements
IT is supported by the Dutch National Science Foundation (NWO Vici VI.C.212.053).
Appendix A Algorithms
@@
The algorithm describing employing the LOWESS method for length bias calibration is shown in Algorithm 1. The main idea is to first fit a locally weighted regression and then adjust the weight based on the residual between regression prediction and the ground truth label. LOWESS is a non-parametric regression technique typically used for smoothing a scatterplot. It is particularly useful when no specific functional form is assumed. Key uses of LOWESS include: (1) Data Smoothing: LOWESS is used to smooth out noise in data, providing a clearer trend by creating a smooth curve that captures the underlying pattern. (2) Data Analysis: By overlaying a LOWESS curve on a scatterplot, one can explore the potential relationship between variables without relying on predefined models or functions, and it’s particularly useful when the relationship between variables is complex and non-linear, as it adapts to the shape of the data. Based on the typical usage of LOWESS, in this paper, we employ LOWESS to investigate the potential relationship between rewards from a biased RM and a characteristic in the output. By assuming that they are actually independent, we regard the prediction from the LOWESS as the bias term and remove it to recover the underlying true reward.
@@
Appendix B Experimental Details
@@
Two key hyper-parameters for our proposed method, RC-Mean and RC-LWR, are the threshold $d$ and the bandwidth $f$. (1) Determine the threshold $d$: The way we use to determine the threshold distance $d$ in Eq. 6 is as follows: (1) Taking the length bias calibration as an example, given a dataset of data pairs to calibrate $\{(x_{1}^{i},x_{2}^{i},r_{\theta}(x_{1}^{i}),r_{\theta}(x_{2}^{i})\}$, we first calculate the average of $||x_{1}^{i}|-|x_{2}^{i}||$, noted as $\operatorname{AVG}(||x_{1}|-|x_{2}||))$, (2) then we set $d=\operatorname{AVG}(||x_{1}|-|x_{2}||))/4$, which is set by intuition. And if the number of data points within the neighborhood is smaller than 10, we do not calibrate that pair as the results may be overly affected by specific data points. This set of hyper-parameters is employed for all three different experimental settings. (2) Determine the bandwidth $f$ for LOWESS: this hyper-parameter is closely related to the assumption. Regarding the experiment of RM as LLM evaluators and RM for LLM alignment, we consider the sufficient density assumption valid since there are 150k samples and 300k samples, respectively. We thus set the bandwidth as default as in statsmodels.api, $f=\frac{1}{3}$, and the number of iterations is 3. (Though we use all scored data points for LOWESS regression, we show in Fig. 12 and Fig. 13 that the methods yield improvements with only hundreds of data points.) Regarding the experiment on the RewardBench, there are only 2,985 data pairs, resulting in 5970 data points. Therefore, the sufficient density assumption may be invalid for some length intervals, so we choose a larger bandwidth $f=0.9$. Importantly, all hyper-parameters relevant to the reward calibration haven’t been tuned for specific reward models. For RewardBench experiments, hyper-parameters for all 33 BT-based reward models and 46 DPO-based reward models are the same, and for AlpacaEval calibration, the hyper-parameters used for 8 models are the same. We argue that this further validates the generality of our proposed methods. In practice, one could conduct a more comprehensive hyper-parameter search for a specific reward model.
@@
Appendix C Extra Experimental Results
@@
Figure 8: Ablation results of different calibration constant on the RewardBench for BT-based reward models.
Table 4: Results of using Length-Calibrated rewards for LLMs alignment with different calibrate constant $\gamma$ in RC-LWR. We apply these methods with four different LLM–RM configurations. The top reports the Length-controlled win rate, and the bottom showcases the average length of generations on AlpacaEval dataset.
@@
Defining the length margin as the absolute difference in length between two completions, the cumulative distribution function of the length margin on four different datasets is shown in Fig. 9. As shown in the Fig. 9, different datasets demonstrate distinct length margin distributions. For Ultrafeedback-Llama3 and Ultrafeedback-Gemma2, about 80% of data pairs have a length margin that is less than 600, which is reasonable because the compared completions are sampled from the same model. And the RewardBench dataset demonstrates a similar pattern. In contrast, the length margin on the AlpacaEval benchmark is a lot longer, with only about 40% of the data points having a length margin that is less than 600. Notably, the simple Length Penalty method works well in calibrated RM on RewardBench and calibrated RM for LLM alignment settings. This may be because the length margin in those two settings is smaller, thus leading to moderate penalization. However, when length margins are more significant, more meticulous adjustments for penalty weight are required as they are more sensitive. In summary, the baseline Length Penalty is already strong when reward models exhibit strong length bias. However, both reward scales and length margins should be considered to adjust the penalty weight, impairing its practicability. In contrast, the proposed method RC-LWR is robust to different hyper-parameter choices, yielding stable improvements in different settings.
@@
Figure 10: Based on the Chat Hard subset of RewardBench, we plot how the performance (left) or improvement brought (right) by RC-LWR vary with the model’s length preference, which is represented by the Spearman score between output length and rewards on RewardBench.
Results in Fig. 4 motivate us to investigate how calibration gains vary for models with different length bias. Therefore, we selected Chat Hard, the most challenging subset of RewardBench, for our investigation. Results are shown in Fig. 10. We use the Spearman correlation to represent the reward models’ length preference on the x-axis; we then plot the model’s performance and RC-LWR improvements, different points in Fig. 10 refers to different models. In Fig. 10 left, we observe that the model’s performance on Char Hard is negatively correlated with the model’s length preference, achieving -0.28 correlation. (Note that a rule-based model that always chooses the shorted sentence could achieve an accuracy of 90% on Chat Hard subset.) In Fig. 10 right, we observe a strong positive correlation between the model’s length preference and its improvement on the Chat Hard subset, achieving 0.76 correlation score. Note that similar trends are found between length preference and overall performance improvement of RewardBench with 0.47 Spearman Correlation score, both indicating that Reward Models with stronger length preference are improved more by the proposed RC-LWR method.

### v2

@@ 6 Conclusion
This paper introduces a post-hoc reward calibration method to mitigate biases in RMs. The proposed approach effectively reduces the impact of prevalent biases, such as length bias, and has been validated across multiple experimental settings. The results consistently show substantial performance gains. Other empirical findings confirm that our method satisfies several critical desiderata for optimal calibration: (1) Our approach consistently mitigates both length and markdown biases in RMs, and can successfully calibrate LLM-as-Judge rewards, underscoring its generality. (2) The method requires no additional data annotation or RM retraining and is computationally efficient. (3) The calibrated rewards exhibit weak correlations with the characteristics of interest, thereby preventing reward hacking and improving the overall quality of the reward signals. (4) For non-biased RMs, our method minimally alters the reward behaviour, whereas it substantially improves RMs with stronger biases, making the calibration impact proportional to the bias level. (5) The method is robust to hyperparameter choices, and a calibration constant can be introduced to preserve some desirable bias towards the specific feature, enhancing its practical usability. (6) The approach is not constrained by dataset size, as even a few hundred data points are sufficient to achieve effective calibration. In conclusion, our approach offers a scalable, efficient, and generalisable solution for addressing biases in RMs, ensuring better alignment between LLMs and human preferences and LLMs evaluation.
Appendix A Length Bias
Length bias emerged as a significant issue in Natural Language Processing (NLP), affecting a wide range of tasks and applications, including machine translation (Zhang et al., 2023; Murray & Chiang, 2018) and text summarization Guo & Vfosoughi (2023); Steen & Markert (2024). This bias occurs when models disproportionately prefer outputs of a certain length, which stands as one of the most consequential sources of bias in NLP.
In LLM era, length bias (also termed as the verbosity issue) continued to be a pressing concern. One key process to develop such LLMs is RLHF, which employ human preference data to fine-tune LLMs (directly or indirectly) to generate more human-aligned outpus but also may unintentionally amplify the length bias in the preference dataset. Therefore, mitigating length bias become innvreasingly important, drawing significant attention. Prior works include: (1) works focusing on RMs training:  Chen et al. (2024b) propose to jointly train two linear reward heads on top of the same model, one trained to correlate with length and the other trained to decorrelate with length, their experiments demonstrate that the approach eliminate the Pearson correlation to 0 (same as our approach); Similarly,  Shen et al. (2023) jointly optimise a main expert RM and a length-bias only expert RM, and exclusively use the main expert in the RL phase; Hou et al. (2024) devise Bucket-Based Length Balancing, that they divide the preference data used to train RM in to different bucket based on length difference, and within each bucket, they balance the number of examples where the better/worse response is more lengthy; (2) works on focusing on LLM alignment:  Meng et al. (2024) leverages the length-normalised reward to prevent model from generating longer but low-quality responses;  Lu et al. (2024) modify the DPO objective by down-sampling same amount of tokens from the chosen and rejected responses when calculating the implicit reward, similar with  Liu et al. (2024b);  Chen et al. (2024a); Liu et al. (2024a); Park et al. (2024b) add an extra penalty term to the implicit reward to penalise the verbosity;  Wu et al. (2024a) set a threshold for top-tier response and select the shortest one as the best response when constructing the preferenc data; and (3) works focusing on the evaluation: Based on AlpacaEval benchmark,  Dubois et al. (2024) propose use LLM ID and instruction ID to do a regression to estimate the bias;  Chiang et al. (2024) employ the same method for Chatbot Arena.
All these previous studies underscore the prevalence and significance of length bias across different subfields of LLMs, highlighting the pressing need for a unified approach to address this issue comprehensively. While our method was not specifically designed to target length bias, it has been empirically validated to effectively mitigate this problem in various contexts, demonstrating its broader applicability.
Appendix B Full Algorithms
@@
The algorithm describing employing the LOWESS method for length bias calibration is shown in Algorithm 1. The main idea is to first fit a locally weighted regression and then adjust the weight based on the residual between regression prediction and the ground truth label. LOWESS is a non-parametric regression technique typically used for smoothing a scatterplot. It is particularly useful when no specific functional form is assumed. Key uses of LOWESS include: (1) Data Smoothing: LOWESS is used to smooth out noise in data, providing a clearer trend by creating a smooth curve that captures the underlying pattern. (2) Data Analysis: By overlaying a LOWESS curve on a scatterplot, one can explore the potential relationship between variables without relying on predefined functions, and it’s particularly useful when the relationship between variables is complex and non-linear, as it adapts to the shape of the data. Based on the typical usage of LOWESS, in this paper, we employ LOWESS to investigate the potential relationship between rewards from a biased RM and a characteristic in the output. By assuming that they are actually independent, we regard the prediction from the LOWESS as the bias term and remove it to recover the underlying true reward.
@@
Appendix C Experimental Details
Experiments with AlpacaEval
AlpacaEval 2.0 is a popular LLM benchmark ranking over 200 models. The official leaderboard employs GPT4-1106-preview as judge. In our setting, we primarily use open-source BT-based RMs to score the LLM’s output. Given a prompt $q$ and corresponding output $y$ from the evaluated LLM and output $y_{b}$ from the baseline LLM, the RM scores them as $r$ and $r_{b}$, respectively; then the Bradley-Terry model yields the preference probability $p(y\succ y_{b}|q)=\sigma(r-r_{b})$. We then compute the average of the perference probability $p(y\succ y_{b}|q)$ across all prompts to determine the final win rate for ranking all evaluated LLMs. We directly use the model’s output from the AlpacaEval official repo. [^3] https://github.com/tatsu-lab/alpaca_eval/tree/main/results We also use our ptoposed method to calibrate AE1. Since AE1 uses GPT4 to compare two responses pairwisely and uses its output logits to compute the preference probability $p(y\succ y_{b}|x)$. Thus, following Eq. 1, we employ the inverse function of $\operatorname{sigmoid}$ to calculate the reward margin $r-r_{b}$. We then utilise the LWR to directly predict the biased reward margin given the length margin $|y|-|y_{b}|$.
@@ Appendix C Experimental Details
Two key hyper-parameters for our proposed method, RC-Mean and RC-LWR, are the threshold $d$ and the bandwidth $f$. (1) Determine the threshold $d$: The way we use to determine the threshold distance $d$ in Eq. 6 is as follows: (1) Taking the length bias calibration as an example, given a dataset of data pairs to calibrate $\{(x_{1}^{i},x_{2}^{i},r_{\theta}(x_{1}^{i}),r_{\theta}(x_{2}^{i})\}$, we first calculate the average of $||x_{1}^{i}|-|x_{2}^{i}||$, noted as $\operatorname{AVG}(||x_{1}|-|x_{2}||))$, (2) then we set $d=\operatorname{AVG}(||x_{1}|-|x_{2}||))/4$, which is set by intuition. And if the number of data points within the neighborhood is smaller than 10, we do not calibrate that pair as the results may be overly affected by specific data points. This set of hyper-parameters is employed for all three different experimental settings. (2) Determine the bandwidth $f$ for LOWESS: this hyper-parameter is closely related to the assumption. Regarding the experiment of RM as LLM evaluators and RM for LLM alignment, we consider the sufficient density assumption valid since there are 150k samples and 300k samples, respectively. We thus set the bandwidth as default as in statsmodels.api, $f=\frac{1}{3}$, and the number of iterations is 3. (Though we use all scored data points for LOWESS regression, we show in Fig. 12 and Fig. 13 that the methods yield improvements with only hundreds of data points.) Regarding the experiment on the RewardBench, there are only 2,985 data pairs, resulting in 5970 data points. Therefore, the sufficient density assumption may be invalid for some length intervals, so we choose a larger bandwidth $f=0.9$. Importantly, all hyper-parameters relevant to the reward calibration haven’t been tuned for specific RM. For RewardBench experiments, hyper-parameters for all 33 BT-based RMs and 46 DPO-based reward models are the same, and for AlpacaEval calibration, the hyper-parameters used for 8 models are the same. We argue that this further validates the generality of our proposed methods. In practice, one could conduct a more comprehensive hyper-parameter search for a specific RM.
@@
Appendix D Extra Experimental Results
@@
Figure 8: Ablation results of different calibration constant on the RewardBench for BT-based RMs.
Table 4: Results of using Length-Calibrated rewards for LLMs’ alignment with different calibrate constant $\gamma$ in RC-LWR. We apply these methods with four different LLM–RM configurations. The top reports the Length-controlled win rate, and the bottom showcases the average length of generations on AlpacaEval dataset.
@@
Defining the length margin as the absolute difference in length between two completions, the cumulative distribution function of the length margin on four different datasets is shown in Fig. 9. As shown in the Fig. 9, different datasets demonstrate distinct length margin distributions. For Ultrafeedback-Llama3 and Ultrafeedback-Gemma2, about 80% of data pairs have a length margin that is less than 600, which is reasonable because the compared completions are sampled from the same LLM. And the RewardBench dataset demonstrates a similar pattern. In contrast, the length margin on the AlpacaEval benchmark is a lot longer, with only about 40% of the data points having a length margin that is less than 600. Notably, the simple Length Penalty method works well in calibrated RM on RewardBench and calibrated RM for LLM alignment settings. This may be because the length margin in those two settings is smaller, thus leading to moderate penalization. However, when length margins are more significant, more meticulous adjustments for penalty weight are required as they are more sensitive. In summary, the baseline Length Penalty is already strong when RMs exhibit strong length bias. However, both reward scales and length margins should be considered to adjust the penalty weight, impairing its practicability. In contrast, the proposed method RC-LWR is robust to different hyper-parameter choices, yielding stable improvements in different settings.
@@
Figure 10: Based on the Chat Hard subset of RewardBench, we plot how the performance (left) or improvement brought (right) by RC-LWR vary with the RM’s length preference, which is represented by the Spearman score between output length and rewards on RewardBench.
Results in Fig. 4 motivate us to investigate how calibration gains vary for RMs with different length bias. Therefore, we selected Chat Hard, the most challenging subset of RewardBench, for our investigation. Results are shown in Fig. 10. We use the Spearman correlation to represent the RMs’ length preference on the x-axis; we then plot the RM’s performance and RC-LWR improvements, different points in Fig. 10 refers to different RMs. In Fig. 10 left, we observe that the RM’s performance on Char Hard negatively correlates with the RM’s length preference, achieving -0.28 correlation. (Note that a rule-based RM that always chooses the shorted sentence could achieve an accuracy of 90% on Chat Hard subset.) In Fig. 10 right, we observe a strong positive correlation between the RM’s length preference and its improvement on the Chat Hard subset, achieving 0.76 correlation score. Note that similar trends are found between length preference and overall performance improvement of RewardBench with 0.47 Spearman Correlation score, both indicating that Reward Models with stronger length preference are improved more by the proposed RC-LWR method.
@@
Further analysis about the chosen length penalty baseline
In our main experiment results, we follow previous research works to set $\alpha=0.001$, which may raise concerns that the $\alpha$ is not optimal for the selected tasks and reward models. Therefore, we first do ablation studies with different $\alpha$ values from $\{0.1,0.05,0.02,0.01,0.005,0.002,0.001,0.0005,0.0002,0.0001\}$. To further decouple the effect of the fact that different reward models have different reward distributions, we propose to first z-normalize the reward and then apply the Length-Penalty method. The results of these two methods are demonstrated in Fig. 15. According to Fig. 15, the chosen value $\alpha=0.001$ achieves the best performance compared with other values (Original line). The Z-normalization trick could slightly boost the performance to 2.75, which is reasonable because it alleviates the issue that different reward models have different distributions. However, both of them underperform the proposed method. Furthermore, we argue that the optimal $\alpha$ is challenging to search for the following reasons: (1) both the reward distribution of the model and the length distribution of the task affect $\alpha$, asking users to search different $\alpha$ per reward model and tasks repeatedly; (2) the method is very sensitive to the value of $\alpha$, for some values the model may even collapse. On the contrary, our proposed method could robustly adapt to different reward models and tasks.
Figure 15: The performance of Length-Penalty baseline with different $\alpha$ values in the RewardbBnech setting. Two different variants are considered here. Original means that we directly add the penalty term to the Reward Model. Normalize means we first z-normalise the rewards and then apply the Length-Penalty baseline.
Reward calibration v.s. Reward Ensemble
As discussed in our related work section, previous works can utilise ensemble techniques to mitigate the potential bias. We thus try to compare our proposed RC-LWR with the ensemble. Specifically, we use normalize-then-average to aggregate rewards from different reward models. We first select the top nine models from the RewardBench Leaderboard as of Aug and then aggregate them. The results are shown in the Fig. 16. In Fig. 16, the number in the cell $[i,j]$ denotes the ensemble performance of model $i$ and model $j$ where $i$ and $j$ are the model’s rank on the RewardBench leaderboard. The last row is the calibrated performance of models. We find that the calibration outperforms most ensemble results except for ensembles with a much stronger model. And intuitively, ensemble-based methods may not address common bias across models. For example, if these models overly prefer lengthy outputs, their ensemble model will also prefer lengthy output.
Figure 16: Ensemble performance and calibrated performance
Visualization
To better illustrate how our proposed method shifts the reward distribution of the model and removes the length bias, we present the kernel density plot (KDE) of rewards (y-axis) w.r.t length (x-axis) for different reward models, calibration methods and hyper-parameters. Specifically, we select ten different reward models (Fig. 17– 26). For each reward mode, we visualize four distributions based on RewardBench dataset: (a) The first (left) figure demonstrates the original distribution between the reward and character length of the example; (b) The second figure demonstrates the calibrated distribution of our proposed RC-LWR method; (c) The third figure demonstrates the calibrated distribution of the proposed variant of RC-LWR (Eq.  8) with $\gamma=1.25$; (d) the last (right) figure illustrates the calibrated distribution of the Length-Penalty baseline with $\alpha=0.001$. In each figure, the colour represents the density of data points. The deeper the colour, the denser the data points. The red dashed line illustrates the LOWESS smoothing for the KDE plot. According to these figures, we find that: (1) As discussed in our introduction, the length bias exists for nearly all selected ten reward models. (Note that the ideal correlation on RewardBench should be slightly negative.) And the correlation is non-linear. (2) The red dashed lines become almost horizontal after RC-LWR calibration, indicating that our method removes the reward’s correlation with length smoothly and uniformly. (3) By setting $\gamma=1.5$, we can push the correlation of rewards and length to be negative, showcasing the potential of our method to inject desirable bias into reward models. (4) Regarding the Length Penalty baseline, though the Spearman correlation between reward and length decreases, the length bias is not removed. For models with no length bias (e.g., NCSOFT/Llama-3-OffsetBias-RM-8B), the length penalty method will push the reward to prefer shorter responses.
Figure 17: internlm/internlm2-20b-reward
Figure 18: NCSOFT/Llama-3-OffsetBias-RM-8B
Figure 19: internlm/internlm2-7b-reward
Figure 20: Ray2333/GRM-llama3-8B-sftreg
Figure 21: Ray2333/GRM-llama3-8B-distill
Figure 22: sfairXC/FsfairX-LLaMA3-RM-v0.1
Figure 23: CIR-AMS/BTRM_Qwen2_7b_0613
Figure 24: openbmb/Eurus-RM-7b
Figure 25: Nexusflow/Starling-RM-34B
Figure 26: internlm/internlm2-1_8b-reward